If I was given the equation 5x-4y=20 what would the slope and y-intercept be. I do not need to graph it.
Thanks

Answer by checkley71(8403)

THE LINE FORMULA IS [Y=mX+b] WHERE m=SLOPE & b=Y INTERCEPT.
5X-4Y=20
-4Y=-5X+20
Y=-5X/-4+20/-4
Y=5X/4-5 HERE YOU HAVE A SLOPE(M)=5/4 & A Y INTERCEPT(b)=-5.